# Seeing Beyond Haze: Generative Nighttime Image Dehazing

### v1

@@ Abstract
Nighttime image dehazing is particularly challenging when dense haze and intense glow severely degrade or completely obscure background information. Existing methods often encounter difficulties due to insufficient background priors and limited generative ability, both essential for handling such conditions. In this paper, we introduce BeyondHaze, a generative nighttime dehazing method that not only significantly reduces haze and glow effects but also infers background information in regions where it may be absent. Our approach is developed on two main ideas: gaining strong background priors by adapting image diffusion models to the nighttime dehazing problem, and enhancing generative ability for haze- and glow-obscured scene areas through guided training. Task-specific nighttime dehazing knowledge is distilled into an image diffusion model in a manner that preserves its capacity to generate clean images. The diffusion model is additionally trained on image pairs designed to improve its ability to generate background details and content that are missing in the input image due to haze effects. Since generative models are susceptible to hallucinations, we develop our framework to allow user control over the generative level, balancing visual realism and factual accuracy. Experiments on real-world images demonstrate that BeyondHaze effectively restores visibility in dense nighttime haze.
Figure 1: Qualitative results from NightDeFog’20 [56], NightEnhance’23 [21], SFSNiD’24 [7] and our method on real-world data. Our method not only reduces dense haze and strong glow effects but also infers missing background details and content in severely degraded regions.
@@ 1 Introduction
The degradation of nighttime imagery due to haze presents a formidable challenge, as the combination of low light conditions and atmospheric particles can significantly obscure scene details. Haze induces scattering and absorption of light, which not only reduces visibility but also creates glows around light sources that conceal the background. The resulting loss of visual information makes it difficult to accurately restore the original scene.
To address this problem, existing supervised nighttime dehazing methods (e.g., [65, 4, 8, 9]) rely primarily on synthetic datasets for training. However, the considerable domain gap between real-world and synthetic degradations—especially in cases of severe haze and glow—limits the effectiveness of their learned background priors. Semi-supervised nighttime dehazing methods (e.g., [7, 21, 56]) combine unlabeled real data with synthetic paired data for training, but their limited background priors and supervision lead to suboptimal performance. In addition, previous methods lack the generative ability to infer background content in regions completely obscured by dense haze or intense glow. As shown in Figure 1, such methods underperform in areas heavily degraded or entirely blocked due to haze.
In this paper, we introduce BeyondHaze, a generative nighttime dehazing method that not only enhances the visibility of scenes affected by haze or glow but also generatively infers background details in regions where the original information is missing. BeyondHaze is built upon two ideas: harnessing the power of image diffusion models to generate clean images by adapting them with dehazing knowledge, and strengthening their generative ability to produce haze-unobscured background content through guided training.
We first pretrain a dehazing model by applying haze and its associated light effects to clear nighttime images and then training the model to restore the original images. This process enables the model to learn robust background priors. After acquiring these priors, the dehazing model generates initial dehazed images and their corresponding confidence maps from real-world haze images. The before-and-after image pairs with confidence maps are then used to fine-tune a pretrained text-to-image diffusion model equipped with LoRA [18], where only the LoRA parameters are updated during training. This approach allows our diffusion model, which possesses strong built-in world priors, to learn from real hazy images and develop generative dehazing capabilities.
@@
Since these training pairs are used for re-creating background information that may be absent in hazy nighttime images, there is a potential to introduce hallucinations. Conversely, omitting these pairs may lead to less visually appealing results but helps to maintain factual accuracy. Our framework is thus designed to give users control over the generative level, allowing them to balance realism and authenticity based on their preferences. We implement this by combining specific text prompts with these training pairs and enabling users to control the generative level of the outputs through prompt selection.
As shown in Figure 1, our method not only reduces dense haze and strong glow effects but also infers missing backgrounds in severely degraded regions. Though the generated details and content may not exactly match those of the actual scene, high-quality generative output nevertheless remains important, as it improves visibility and overall image quality, especially in regions where precise background recovery is not possible. This approach yields more visually realistic predictions, which is particularly valuable for image editing applications. Moreover, by adjusting text prompts, we can control the level of generative inference.
Experimental results on real-world datasets demonstrate that BeyondHaze achieves significant performance improvements. Notably, our method achieves a MUSIQ score of 65.79 and a ClipIQA score of 0.6774, surpassing existing nighttime image dehazing methods by 27.4% in MUSIQ and 8.9% in ClipIQA. The main contributions of this work are as follows:
@@
  • The generative ability of the diffusion model is improved for nighttime dehazing via training pairs produced by our detail enhancement model and severe degradation model.
  • Control over the generative level of the output is made possible through the association of specific text prompts with the training pairs.
  • Both qualitative and quantitative results demonstrate that BeyondHaze significantly improves the quality of real-world nighttime haze images. Experimental results show that it not only suppresses degradation factors like haze and glow but also enhances overall image quality.
Figure 2: Our approach integrates dehazing priors with generative capabilities to enhance nighttime image dehazing. We first pretrain a dehazing model using augmentations that simulate noise and light effects in hazy night scenes. The knowledge from this dehazing model is then distilled into a pretrained image diffusion model with LoRA. Additional training pairs are generated by two supplementary models: a detail enhancement model that applies super-resolution to initial dehazed images, and a severe degradation model which produces substantially obscured images from real-world clear images. These training pairs, combined with customized text prompts, are used to fine-tune the diffusion model and enable controllable generative dehazing.
@@ 2 Related Work
Traditional image dehazing methods [49, 11, 15, 3, 37] address haze effects by estimating transmission maps and atmospheric light. However, these methods struggle with nighttime haze images, as the lower signal-to-noise ratio of nighttime images significantly affects the accuracy of transmission map and atmospheric light estimation. Existing daytime image dehazing methods [5, 43, 62, 25, 44, 29, 42, 59, 60, 35, 10, 41, 67, 55, 14, 47, 28, 46, 6, 30, 26, 27, 19, 13, 34, 66, 58] can be retrained on nighttime haze datasets. However, current nighttime haze datasets are synthetic, and models trained on them struggle to handle real-world haze due to the significant domain gap between real and synthetic degradations.
Existing nighttime image dehazing methods can be roughly divided into two categories: optimization-based and learning-based. The optimization-based dehazing methods [40, 1, 2, 63, 64, 50, 36, 52] utilize the atmospheric scattering model or a novel imaging model to reduce nighttime haze effects. However, nighttime haze images are also influenced by other complex factors, such as noise and glow effects, which significantly affect the accuracy of these models.
Learning-based nighttime image dehazing methods [65, 56, 24, 21] utilize paired synthetic data or unpaired real-world data to train neural networks. However, there is a significant domain gap between real-world and synthetic degradations that lead to inaccuracies when handling real-world haze images. Unsupervised and semi-supervised nighttime image dehazing methods [56, 21, 31, 7] primarily use unpaired data or pseudo ground truths for training. However, unpaired data lacks accurate pixel-level supervision, and pseudo ground truths often do not accurately represent areas with severe degradation due to the limitations of teacher models. Furthermore, these methods face challenges in recovering substantially obscured regions due to haze or low-light conditions because the true signal in these areas is much smaller than the degradation. To mitigate this issue, we utilize large-scale generative models, such as text-to-image diffusion models, to enhance visibility in nighttime haze images. By leveraging the strong generative capabilities of these models, our method is better equipped to handle haze effects, particularly in dense haze regions and extreme low-light conditions.
@@
Diffusion models [17] have achieved remarkable success in various image restoration tasks, including deraining [33, 32], desnowing [38], and removing raindrops [22]. However, their generalization ability is often constrained by the quality of training datasets. For nighttime image dehazing, the lack of real-world paired training data poses a significant challenge. Our work addresses this issue by utilizing different types of training data that target different aspects of nighttime dehazing, where each aspect is treated using reliable data for the corresponding sub-problem (i.e., training pairs for dehazing without recovery of missing background information from a state-of-the-art dehazing model trained on real nighttime haze images, training pairs for detail enhancement by a pre-trained super-resolution model applied to dehazed images, and training pairs for missing background generation by a severe degradation model that obscures real-world image regions that would be heavily impacted by haze).
@@ 3 Proposed Method: BeyondHaze
Figure 2 illustrates the overall pipeline of BeyondHaze, which is based on generative dehazing priors and controllable generative dehazing. The generative dehazing priors are learned by distilling the knowledge of a pre-trained dehazing model into an image diffusion model with LoRA. Background detail and content generation are learned through training data from a detail enhancement model and a severe degradation model, where the generative level of the output is regulated by text prompts.
@@ 3.1 Generative Dehazing Priors
Our pre-trained dehazing model is a vision transformer [16] trained with severe augmentations [31], where clear images are intentionally degraded with light effects and noise. This forces the model to learn robust background priors, enabling it to infer clear scene information despite the degraded input.
Once the pre-trained dehazing model is obtained, we use it to generate initial dehazed images from real-world hazy nighttime images. These image pairs are then used to train our diffusion model, enabling it to progressively acquire task-specific dehazing knowledge while preserving its inherent generative properties. We refer to these integrated priors within the diffusion model as generative dehazing priors. Optimizing a large-scale diffusion model is, however, computationally expensive and risks degrading its generative properties. To address this, we employ LoRA [18], updating only its parameters during training. This preserves the model’s generative properties while incorporating dehazing capabilities.
@@ 3.2 Controllable Generative Dehazing
While the training data from our detail enhancement model can help the diffusion model learn to synthesize fine-scale details missing from the initial dehazed images, it does not help with recovering background regions occluded by dense haze or strong glow. To address this, we introduce a severe degradation model to produce training data for improving such generative capabilities in diffusion models. The degradation model severely degrades clear nighttime images by blending light effects and adding noise, creating a heavily augmented dataset, formulated as in [31]:
@@
Metrics Uformer Restormer DiT NightDeFog NightEnhance SFSNiD Ours Ours
CVPR’22 CVPR’22 ICCV’23 ECCV’20 ACM MM’23 CVPR’24 w/o Gen
MUSIQ 44.33 45.04 46.02 51.61 50.56 43.28 60.29 65.79
TRES 51.58 51.89 53.87 68.88 59.08 38.50 68.05 80.08
ClipIQA 0.3256 0.3762 0.4316 0.3743 0.3673 0.3672 0.6217 0.6774
HyperIQA 0.4035 0.4068 0.4122 0.4462 0.3981 0.2839 0.4831 0.6139
ManIQA 0.3050 0.3210 0.3311 0.3273 0.2879 0.2468 0.3690 0.4964
Table 1: Quantitative comparison on the RealNightHaze dataset (evaluated at 512$\times$512 resolution). Five non-reference metrics, namely MUSIQ [23], TRES [12], ClipIQA [51], HyperIQA [48], and ManIQA [57], are used for evaluation. For all five metrics, higher values are better. NightDeFog [56], NightEnhance [21] and SFSNiD [7] are nighttime image dehazing methods, while the remaining methods are image restoration backbones. “w/o Gen” denotes that neither generative detail enhancement nor region synthesis are selected.

### v2

@@ Abstract
Nighttime image dehazing is particularly challenging when dense haze and intense glow severely degrade or entirely obscure background information. Existing methods often struggle due to insufficient background priors and limited generative capability, both of which are highly important under such conditions. In this paper, we introduce BeyondHaze, a generative nighttime dehazing method that not only reduces haze and glow effects but also reconstructs plausible background structures in regions where visual cues are heavily degraded. Our approach is built on two main ideas: obtaining strong background priors by adapting image diffusion models to nighttime dehazing, and enhancing generative ability in haze- and glow-obscured areas through guided training. Task-specific nighttime dehazing knowledge is distilled into an image diffusion model while preserving its capacity to generate clean images. The diffusion model is further trained on tailored image pairs to improve its ability to recover background details that are suppressed by haze effects. Since generative models may introduce hallucinated content, we design our framework to allow user control over the generative level, enabling a balance between visual realism and fidelity. Experiments on real-world nighttime images demonstrate that BeyondHaze substantially improves visibility and scene detail under dense haze.
Figure 1: Qualitative results from NightDeFog’20 [yan2020nighttime], NightEnhance’23 [jin2023enhancing], SFSNiD’24 [cong2024semi] and our method on real-world data. Our method not only reduces dense haze and strong glow but also infers missing background details and content in severely degraded regions.
@@ 1 Introduction
The degradation of nighttime imagery due to haze presents a formidable challenge, as the combination of low light conditions and atmospheric particles can significantly obscure scene details. Haze induces scattering of light, which not only reduces visibility but also creates glows around light sources that conceal the background. The resulting loss of visual information makes it difficult to accurately restore the original scene.
To address this problem, existing supervised nighttime dehazing methods (e.g., [zhang2020nighttime, bernabel2024ndels, cui2023focal, cui2024omni]) rely on synthetic datasets for training. However, the considerable domain gap between real-world and synthetic degradations, especially in cases of severe haze and glow, limits the effectiveness of their learned background priors. Semi-supervised nighttime dehazing methods (e.g., [cong2024semi, jin2023enhancing, yan2020nighttime]) combine unlabeled real data with synthetic paired data for training, but their reliance on pseudo-labels and weak supervision constrains their performance. In addition, previous methods lack the generative ability to infer background content in regions completely obscured by dense haze or intense glow. As shown in Figure 5, such methods underperform in areas heavily degraded or entirely blocked due to haze.
In this paper, we introduce BeyondHaze, a generative nighttime dehazing method that not only enhances the visibility of scenes affected by haze or glow but also generates plausible background details in regions where the original information is missing. BeyondHaze is built upon two ideas: adapting image diffusion models with nighttime dehazing knowledge to enable clean-image generation, and strengthening their generative capability to produce background content that is absent due to severe haze through guided training.
We first pretrain a dehazing model by applying haze and its associated light effects to clear nighttime images and then training the model to restore the original images. This process enables the model to learn robust background priors. After acquiring these priors, the dehazing model generates initial dehazed images and their corresponding confidence maps from real haze images. The before-and-after image pairs with confidence maps are then used to fine-tune a pretrained diffusion model equipped with LoRA [hu2021lora]. This approach allows our diffusion model, which possesses strong built-in world priors, to learn from real haze images and develop generative dehazing capabilities.
@@
Using these training pairs enables the model to recreate missing background information, but also introduces the risk of hallucination. Without them, results are more faithful but less visually appealing. To balance these trade-offs, our framework allows users to control the generative level through text prompts. During inference, we also compute a generative-level map for the high-generative output, indicating how strongly each region relies on generative synthesis rather than factual restoration. This map is derived from the attention scores between tokens obtained from the high- and low-generative inference paths.
As shown in Figure 5, our method not only reduces dense haze and strong glow but also recovers plausible background content in regions where visibility is severely degraded. In areas where the scene is fully obscured, exact recovery is unattainable, yet generating a visually coherent estimate remains useful, especially for image editing tasks. Importantly, users can control the generative level through text prompts, allowing the output to range from conservative restoration to stronger generative completion. This controllable design enables users to balance faithful recovery and visual completeness depending on the application.
Experimental results on real-world datasets show that BeyondHaze achieves significant performance improvements. Notably, our method achieves a MUSIQ score of 65.79 and a ClipIQA score of 0.6774, surpassing existing nighttime image dehazing methods by 27.4% in MUSIQ and 8.9% in ClipIQA. As a summary:
@@
  • Control over the generative level of the output is made possible through the association of specific text prompts with the training pairs. A generative-level map is further derived during inference to reveal potential hallucination-prone regions in the outputs.
  • Both qualitative and quantitative results show that BeyondHaze significantly improves real-world nighttime haze images, effectively suppressing haze and glow while enhancing overall image quality.
@@ 2 Related Work
Existing daytime image dehazing methods [cai2016dehazenet, ren2016single, zhang2018densely, li2017aod, ren2018gated, li2018single, qu2019enhanced, ye2022perceiving, yu2022frequency, liu2019griddehazenet, dong2020multi, qin2020ffa, zheng2021ultra, wu2021contrastive, guo2022image, song2023vision, li2019semi, shao2020domain, chen2021psd, li2022physically, li2020zero, li2021you, huang2019towards, golts2019unsupervised, liu2020end, zhao2021refinednet, yang2022self, chen2024dual] can be retrained on nighttime haze datasets. However, current nighttime haze datasets are synthetic, and models trained on them struggle to handle real-world haze due to the significant domain gap between real and synthetic degradations.
Existing nighttime image dehazing methods can be roughly divided into two categories: optimization-based and learning-based. The optimization-based dehazing methods [pei2012nighttime, ancuti2016night, ancuti2020day, zhang2014nighttime, zhang2017fast, tang2021nighttime, liu2022nighttime, wang2022variational] utilize the atmospheric scattering model or a novel imaging model to reduce nighttime haze effects. However, nighttime haze images are also influenced by other complex factors such as noise and glow, which significantly affect the accuracy of these models.
Learning-based nighttime dehazing methods [zhang2020nighttime, kuanar2022multi, jin2023enhancing] rely on paired synthetic or unpaired real data, but the domain gap between synthetic and real degradations leads to inaccuracies on real haze images. Unsupervised and semi-supervised approaches [jin2023enhancing, lin2024nighthaze, cong2024semi] use unpaired data or pseudo ground truths, yet unpaired training lacks pixel-level supervision and pseudo labels often fail in heavily degraded regions due to teacher-model limitations. Moreover, these methods struggle to recover regions where the true scene signal is overwhelmed by by severe haze or noise.
@@
Diffusion models [ho2020denoising] have shown strong performance in image restoration tasks [lin2025rgb, liu2024residual, ozdenizci2023restoring, lin2025geocomplete] such as deraining [lin2024nightrain] and raindrop removal [jin2024raindrop]. Their effectiveness, however, depends heavily on training data quality, which is a major limitation for nighttime dehazing due to the lack of real paired datasets. Recent approaches [lan2025exploiting, tang2024realfill, parmar2024one] leverage the generative priors of pretrained diffusion models, but supervised methods [tang2024realfill, parmar2024one] remain constrained by dataset fidelity, and unpaired frameworks [lan2025exploiting] lack precise pixel-level supervision and struggle to learn strong background priors.
Our work tackles these challenges by leveraging multiple types of training data, each tailored to a specific sub-problem of nighttime dehazing. We use (1) dehazing pairs obtained by applying a SOTA dehazing model to real nighttime haze images to create pseudo ground truths without hallucinating missing backgrounds, (2) enhancement pairs generated by applying a pretrained super-resolution model to dehazed images, and (3) generative pairs produced by heavily degrading real images to simulate regions where background content would be obscured by haze.
Figure 2: Our approach integrates dehazing priors with generative capabilities to enhance nighttime image dehazing. We first pretrain a dehazing model using augmentations that simulate noise and light effects in hazy night scenes. The knowledge from this dehazing model is then distilled into a pretrained image diffusion model with LoRA. Additional training pairs are generated by two supplementary models: a detail enhancement model that applies super-resolution to initial dehazed images, and a severe degradation model which produces substantially obscured images from clear images. These training pairs, combined with customized text prompts, are used to fine-tune the diffusion model and enable controllable generative dehazing. During inference, the fine-tuned model performs two inference branches conditioned on low- and high-generative text prompts, producing the corresponding outputs. A generative-level map is then derived by computing attention scores between the two branches, revealing regions that rely more on generative synthesis than on factual restoration.
@@ 3 Proposed Method: BeyondHaze
Figure 2 illustrates the overall pipeline of BeyondHaze, which is based on generative dehazing priors and controllable generative dehazing. The generative dehazing priors are learned by distilling the knowledge of a pre-trained dehazing model into an image diffusion model with LoRA. Background detail and content generation are learned through training data from a detail enhancement model and a severe degradation model, where the generative level of the output is regulated by text prompts. During inference, the generative level of the outputs can be controlled by adjusting the text prompts. Furthermore, a generative-level map is obtained by calculating attention scores between tokens from the low- and high-generative inference paths, revealing potential hallucination-prone regions.
@@ 3.1 Generative Dehazing Priors
Our pre-trained dehazing model is a vision transformer [he2022masked] trained with severe augmentations [lin2024nighthaze], where clear images are intentionally degraded with light effects and noise. This forces the model to learn robust background priors, enabling it to infer clear scene information despite degraded inputs.
Once the pre-trained dehazing model is obtained, we use it to generate initial dehazed images from real-world hazy nighttime images. These image pairs are then used to train our diffusion model, enabling it to progressively acquire task-specific dehazing knowledge while preserving its inherent generative properties. We refer to these integrated priors within the diffusion model as generative dehazing priors. Optimizing a large-scale diffusion model is, however, computationally expensive and risks degrading its generative properties. To address this, we employ LoRA [hu2021lora], updating only its parameters during training. This preserves the model’s generative properties while incorporating dehazing capabilities.
@@ 3.2 Controllable Generative Dehazing
While the training data from our detail enhancement model can help the diffusion model learn to synthesize fine-scale details missing from the initial dehazed images, it does not help with recovering background regions occluded by dense haze or strong glow. To address this, we introduce a severe degradation model to produce training data for improving such generative capabilities in diffusion models. The degradation model severely degrades clear nighttime images by blending light effects and adding noise, creating a heavily augmented dataset, formulated as in [lin2024nighthaze]:
@@
3.3 Inference with Generative-Level Map
During inference, our fine-tuned diffusion model can produce dehazed results with varying generative levels by using controllable text prompts. Specifically, we define two types of prompts: a low-generative prompt $t_{\text{low}}$, which prioritizes faithful restoration with higher factual accuracy but lower perceptual richness, and a high-generative prompt $t_{\text{high}}$, which encourages the synthesis of missing details and visually enhanced content. Given an input hazy image $\mathbf{x}$, the model performs two inference branches conditioned on these prompts, producing two corresponding outputs that represent the low- and high-generative results, respectively.
To quantify how much each region in the generative output relies on synthesis, we analyze the relationship between tokens from the two branches. In the $\ell$-th attention layer, the tokens from the high- and low-generative branches are concatenated along the token dimension and then projected into query and key embeddings. We define the merged representations as $\mathbf{Q}_{\text{merge}}^{(\ell)}=[\mathbf{Q}_{\text{high}}^{(\ell)};\mathbf{Q}_{\text{low}}^{(\ell)}]$ and $\mathbf{K}_{\text{merge}}^{(\ell)}=[\mathbf{K}_{\text{high}}^{(\ell)};\mathbf{K}_{\text{low}}^{(\ell)}]$, where $\cdot$$\cdot$] denotes concatenation along the token dimension. The attention is then computed as
$\mathbf{A}^{(\ell)}=\mathrm{Softmax}\!\left(\frac{\mathbf{Q}_{\text{merge}}^{(\ell)}\left(\mathbf{K}_{\text{merge}}^{(\ell)}\right)^{\top}}{\sqrt{d}}\right),$ (8)
where $d$ is the feature dimension.
Among the merged attention matrix $\mathbf{A}^{(\ell)}\in\mathbb{R}^{2N\times 2N}$, where $N$ denotes the number of tokens in each branch, the submatrix corresponding to the high-generative queries $\mathbf{Q}_{\text{high}}^{(\ell)}\in\mathbb{R}^{N\times d}$ can be written as
$\mathbf{A}^{(\ell)}_{\text{high}}=[\mathbf{A}^{(\ell)}(\mathbf{Q}_{\text{high}}^{(\ell)},\mathbf{K}_{\text{high}}^{(\ell)})\;;\;\mathbf{A}^{(\ell)}(\mathbf{Q}_{\text{high}}^{(\ell)},\mathbf{K}_{\text{low}}^{(\ell)})]\in\mathbb{R}^{N\times 2N},$
where each row represents the attention probability distribution of a high-generative token over the high- and low-generative keys. These probabilities indicate how likely each token relies on generative or restorative information. A higher probability toward $\mathbf{K}_{\text{high}}^{(\ell)}$ and a lower probability toward $\mathbf{K}_{\text{low}}^{(\ell)}$ suggest that the region is more generative, while the opposite indicates a more restoration-dominated area.
The final generative and restoration scores are obtained by averaging the layer-wise attention probabilities across all $L$ layers:
$s_{\text{gen}}$ $=$ $\frac{1}{L}\sum_{\ell=1}^{L}\sum_{q\in\mathbf{K}_{\text{high}}^{(\ell)}}\mathbf{A}^{(\ell)}(\mathbf{Q}_{\text{high}}^{(\ell)},q),$ (9)
$s_{\text{res}}$ $=$ $\frac{1}{L}\sum_{\ell=1}^{L}\sum_{q\in\mathbf{K}_{\text{low}}^{(\ell)}}\mathbf{A}^{(\ell)}(\mathbf{Q}_{\text{high}}^{(\ell)},q).$ (10)
where $s_{\text{gen}},s_{\text{res}}\in\mathbb{R}^{N\times 1}$ represent the averaged probabilities of attending to the high- and low-generative keys, respectively. For visualization, these scores are resized to the original image resolution to obtain the final generative-level map, which reflects the spatial distribution of generative strength across the image. The attention computation for the generative-level map is performed solely for score extraction and does not affect the actual inference results. To ensure that the two inference branches remain independent, an attention mask is applied to prevent any information exchange between the low- and high-generative tokens.
Table 1: Quantitative comparison on the RealNightHaze dataset (evaluated at $\times512$ resolution). Five non-reference metrics, namely MUSIQ [ke2021musiq], TRES [golestaneh2022no], ClipIQA [wang2023exploring], HyperIQA [su2020blindly], and ManIQA [yang2022maniqa], are used for evaluation. For all five metrics, higher values are better. NightDeFog [yan2020nighttime], NightEnhance [jin2023enhancing], Diff-Dehazer [lan2025exploiting] and SFSNiD [cong2024semi] are image dehazing methods, while the remaining methods are image restoration backbones. “w/o Gen” denotes that neither generative detail enhancement nor region synthesis are selected.
Metrics Uformer Restormer DiT RealFill NightDeFog NightEnhance Diff-Dehazer SFSNiD Ours w/o Gen Ours
MUSIQ $\uparrow$ 44.33 45.04 46.02 51.73 51.61 50.56 44.57 43.28 60.29 65.79
TRES $\uparrow$ 51.58 51.89 53.87 50.25 68.88 59.08 50.59 38.50 68.05 80.08
ClipIQA $\uparrow$ 0.3256 0.3762 0.4316 0.4771 0.3743 0.3673 0.4136 0.3672 0.6217 0.6774
HyperIQA $\uparrow$ 0.4035 0.4068 0.4122 0.3556 0.4462 0.3981 0.3939 0.2839 0.4831 0.6139
ManIQA $\uparrow$ 0.3050 0.3210 0.3311 0.2656 0.3273 0.2879 0.3055 0.2468 0.3690 0.4964